OpenStudy (anonymous):

The equation y = x - 3 represents a line perpendicular to y = -x + 3. Is This TRUE or FALSE?

OpenStudy (anonymous):

finding somethign perpendicular means teh number before the x is opposite resiprical. the first equation has 1. and the second is -1. TRUE

OpenStudy (anonymous):

y1 = x - 3 so m1 = 1 y2 = -x + 3 so m1 = -1 for the two lines to be perpendicular it must satisfy m1 = -1/m2 since 1 = -1/(-1) then yes in fact they are perpendicular. TRUE.

OpenStudy (anonymous):

m2 = -1 ****